OpenStudy (anonymous):

Let me try this again the problem reads: X squared -3X-5 for X=-6

OpenStudy (aroub):

This is your question : x^2-3x-5 and x=-6 ?

OpenStudy (anonymous):

no sorry it's x^-3x-5 and x=-6

OpenStudy (aroub):

x^-3x-5 and x=-6 All you have to do is substitute x by -6 => -6^-3(-6)-5 -6^(18)-5 -101559956668416-5 -101559956668421 I think this is the way you do it. The answer is pretty long :)

OpenStudy (aroub):

No wait..

OpenStudy (aroub):

It depends on how the question is if it's like this : x^(-3x)-5 then the answer is -101559956668421 but if it was like this : x^(-3x-5) => -6^(-3(-6)-5) -13060694016

OpenStudy (phi):

@dion Is this multiple choice? If so, can you post the choices? I am guessing that the problem really is x^2 - 3x -5, evaluated at x=-6 . As you see form Aroub's answers, for problems with x^x, the answers are very large. to figure these problems out, you need to know that x^2 means x times x, 3x means 3 times x, etc. you replace the x with -6 in the formula: (-6)*(-6) - 3* (-6) -5 now it is a problem with multiplying and adding negative/positive numbers. -6*-6= +36 (two minus signs change to 1 plus sign) 3*-6= -18 so you have +36 - -18 -5 - -18 (subtracting a negative 18) means add (change 2 minus signs to a plus) 36+18-5= 49